OpenStudy (anonymous):

Find S8 for the series -2+-10+-50+-250..

OpenStudy (anonymous):

sn = a1(1 - r^n) / 1 - r n = 8 a1 = -2 r = 5 s8 = -2(1 - 5^8)/(1 - 5) s8 = -2(1 - 390625)/-4 s8 = (-2 * -390624)/-4 s8 = 781248/-4 s8 = -195312 you might need a second opinion....

OpenStudy (anonymous):

You were right! Thank you so much!^^

OpenStudy (anonymous):

your welcome :)